OpenStudy (anonymous):

solve for m: s= sqare root of 8mn

OpenStudy (anonymous):

\[m = (s^2) / 8n\]

OpenStudy (anonymous):

Can you please show me the steps

OpenStudy (anonymous):

Yeah. Square both sides: \[s^2 = 8mn\] Divide both sides by 8: \[s^2/8 = mn\] Divide both sides by n: \[s^2/8n = m\]